Problem 9-14 Treasury STRIPS (LO3, CFA1)

U.S. Treasury STRIPS, close of business November 15, 2015:

Maturity Price Maturity Price
November ’16 99.471 November ’19 95.035
November ’17 98.782 November ’20 92.570
November ’18 96.827 November ’21 89.342

a. Calculate the quoted yield for each of the STRIPS given in the table above. (Input all amounts as positive values. Do not round intermediate calculations. Enter your answers as a percent rounded to 3 decimal places.)

b. Does the market expect interest rates to go up or down in the future?
Answer #1

a). Price of STRIP = Face Value / (1 + r)n

STRIPS are semi-annually compounded.

For November '16:

99.471 = 100 / [1 + (y/2)]2

[1 + (y/2)]2 = 100 / 99.471

1 + (y/2) = [1.0053]1/2

y/2 = 1.0027 - 1

y = 0.0027 * 2 = 0.0053, or 0.53%

For November '17:

98.782 = 100 / [1 + (y/2)]4

[1 + (y/2)]4 = 100 / 98.782

1 + (y/2) = [1.0123]1/4

y/2 = 1.0031 - 1

y = 0.0031 * 2 = 0.0061, or 0.61%

For November '18:

96.827 = 100 / [1 + (y/2)]6

[1 + (y/2)]6 = 100 / 96.827

1 + (y/2) = [1.0328]1/6

y/2 = 1.0054 - 1

y = 0.0054 * 2 = 0.0108, or 1.08%

For November '19:

95.035 = 100 / [1 + (y/2)]8

[1 + (y/2)]8 = 100 / 95.035

1 + (y/2) = [1.0522]1/8

y/2 = 1.0064 - 1

y = 0.0064 * 2 = 0.0128, or 1.28%

For November '20:

92.570 = 100 / [1 + (y/2)]10

[1 + (y/2)]10 = 100 / 92.570

1 + (y/2) = [1.0803]1/10

y/2 = 1.00775 - 1

y = 0.00775 * 2 = 0.0155, or 1.55%

For November '21:

89.342 = 100 / [1 + (y/2)]12

[1 + (y/2)]12 = 100 / 89.342

1 + (y/2) = [1.1193]1/12

y/2 = 1.0094 - 1

y = 0.0094 * 2 = 0.0189, or 1.89%

b). Note that the term structure is upward sloping; the expectations hypothesis then implies that this reflects market expectations of rising interest rates in the future.